# Diffusion Model-Based Image Editing: A Survey

### v2

@@ 3 Categorization of Image Editing
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=2em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=4.5em,font=,, where level=2text width=3.1em,font=,, where level=3text width=3.0em,font=,, [ Training-Based
Approaches [ Domain-Specific Editing
with Weak Supervision, text width=7.5em [ DiffusionCLIP [131], Asyrp [132], EffDiff [133], DiffStyler [134], StyleDiffusion [135],
UNIT-DDPM [136], CycleNet [137], Diffusion Autoencoders [138], HDAE [139],
EGSDE [140], Pixel-Guided Diffusion [141] , leaf, text width=27em ] ] [ Reference and Attribute
Guidance via
Self-Supervision, text width=7.5em [ PbE [142], RIC [143], ObjectStitch [144], PhD [145], DreamInpainter [146],
Anydoor [147], FADING [148], PAIR Diffusion [149], SmartBrush [150],
IIR-Net [151], PowerPaint [152], Imagen Editor [153], SmartMask [154],
Uni-paint [155] , leaf, text width=24em ] ] [ Instructional Editing
via Full Supervision, text width=6.5em [ InstructPix2Pix [156], MoEController [157], FoI [158], LOFIE [159],
InstructDiffusion [160], Emu Edit [161], DialogPaint [162], Inst-Inpaint [163],
HIVE [164], ImageBrush [165], InstructAny2Pix [166], MGIE [167], SmartEdit [168] , leaf, text width=27em ] ] [ Pseudo-Target Retrieval
with Weak Supervision, text width=7.5em [ iEdit [169], TDIELR [170], ChatFace [171] , leaf, text width=14em ] ] ]
@@
In addition to the significant advancements that diffusion models have achieved in image generation, restoration, and enhancement, they also make notable breakthroughs in image editing by offering enhanced controllability compared to the previously dominant GANs. Unlike image generation that focuses on creating new images from scratch, and image restoration and enhancement that aims repairing and improving the quality of degraded images, image editing involves the modification of existing images in the aspects of appearance, structure, or content, including tasks like adding objects, replacing backgrounds, and altering textures.
In this survey, we organize image editing papers into three principle groups based on their learning strategies: training-based approaches, testing-time finetuning approaches, and training and finetuning free approaches, which are elaborated in Sections 45, and 6, respectively. Additionally, we explore 10 types of input conditions employed by these methods to control the editing process, including text, mask, reference (Ref.) image, class, layout, pose, sketch, segmentation (Seg.) map, audio, and dragging points. Furthermore, we investigate 12 most common editing types that can be accomplished by these methods, which are organized into three broad categories defined as follows.
@@ 4 Training-Based Approaches
In the field of diffusion model-based image editing, training-based approaches have gained significant prominence. These approaches are not only notable for their stable training of diffusion models and their effective modeling of data distribution but also for their reliable performance in a variety of editing tasks. To thoroughly examine these methods, we categorize them into four main groups based on their application scopes, the conditions required for training, and the types of supervision, as shown in Fig. 2. Further, within each of these primary groups, we classify the approaches into distinct types based on their core editing methodologies. This classification illustrates the range of these methods, from targeted domain-specific applications to broader open-world uses.
4.1 Domain-Specific Editing with Weak Supervision
Over the past several years, Generative Adversarial Networks (GANs) have been widely explored in image editing for their ability to generate high-quality images. However, diffusion models, with their advanced image generation capabilities, emerge as a new focus in this field. One challenge with diffusion models is their need for extensive computational resources when trained on large datasets. To address this, earlier studies train these models via weak supervision on smaller specialized datasets. These datasets highly concentrate on specific domains such as CelebA [236] and FFHQ [2] for human face manipulation, AFHQ [237] for animal face editing and translation, LSUN [238] for object modification, and WikiArt [239] for style transfer. To thoroughly understand these approaches, we organize them according to their types of weak supervision.
(a) Simplified training pipeline of DiffusionCLIP [131]
(b) Simplified training pipeline of Asyrp [132]
Figure 3: Comparison of the training pipelines between two representative CLIP guided methods: DiffusionCLIP [131] and Asyrp [132]. The sample images are from DiffusionCLIP [131] and Asyrp [132] on the AFHQ [237] dataset.
CLIP Guidance. Drawing inspiration from GAN-based methods [240, 241] that use CLIP [242] for guiding image editing with text, several studies incorporate CLIP into diffusion models. A key example is DiffusionCLIP [131], which allows for image manipulation in both trained and new domains using CLIP. Specifically, it first converts a real image into latent noise using DDIM inversion, and then finetunes the pretrained diffusion model during the reverse diffusion process to adjust the image’s attributes constrained by a CLIP loss between the source and target text prompts. Instead of finetuning on the whole diffusion model, Asyrp [132] focuses on a semantic latent space internally, termed h-space, where it defines an additional implicit function parameterized by a small neural network. Then it trains the network with the guidance of the CLIP loss while keeping the diffusion model frozen. A visual comparison of the simplified pipelines of DiffusionCLIP and Asyrp is shown in Fig. 3. To address the time-consuming problem of multi-step optimization in DiffusionCLIP, EffDiff [133] introduces a faster method with single-step training and efficient processing.
@@ 4.1 Domain-Specific Editing with Weak Supervision
Cycling Regularization. Since diffusion models are capable of domain translation, the cycling framework commonly used in methods like CycleGAN [5] is also explored within them. For instance, UNIT-DDPM [136] defines a dual-domain Markov chain within diffusion models, using cycle consistency to regularise training for unpaired image-to-image translation. Similarly, CycleNet [137] adopts ControlNet [101] with pretrained Stable Diffusion [26] as the backbone for text conditioning. It also employs consistency regularization throughout the image translation cycle, which encompasses both a forward translation from the source domain to the target domain and a backward translation in the reverse direction.
@@
4.2 Reference and Attribute Guidance via Self-        Supervision
@@
4.3 Instructional Editing via Full Supervision
@@
4.4 Pseudo-Target Retrieval with Weak Supervision
@@ 5 Testing-Time Finetuning Approaches
In image generation and editing, testing-time finetuning represents a significant step towards precision and control. This section explores various finetuning strategies (Fig. 5) that enhance image editing capabilities. These approaches, as shown in Fig.  6, range from finetuning the entire denoising model to focusing on specific layers or embeddings. We investigate methods that finetune the entire model, target specific parameters, and optimize text-based embeddings. Additionally, we discuss the integration of hypernetworks and direct image representation optimization. These methods collectively demonstrate the evolving complexity and effectiveness of finetuning techniques in image editing, catering to a wide array of editing requirements and user intentions.
@@ 5.1 Denoising Model Finetuning
Finetuning Entire Denoising Models. Finetuning the entire denoising model allows the model to better learn specific features of images and more accurately interpret textual prompts, resulting in edits that more closely align with user intent. UniTune [172] finetunes the diffusion model on a single base image during the tuning phase, encouraging the model to produce images similar to the base image. During the sampling phase, a modified sampling process is used to balance fidelity to the base image and alignment to the editing prompt. This includes starting sampling from a noisy version of the base image and applying classifier-free guidance during the sampling process. Custom-Edit [173] uses a small set of reference images to customize the diffusion model, enhancing the similarity of the edited image to the reference images while maintaining the similarity to the source image.
@@
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=2em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=4.5em,font=,, where level=2text width=3.1em,font=,, where level=3text width=3.0em,font=,, [ Testing-Time
Finetuning Approaches [ Denosing Model
Finetuning, text width=5.5em [ UniTune [172], Custom-Edit [173], KV-Inversion [174] , leaf, text width=17em ] ] [ Embedding
Finetuning, text width=4em [ Null-Text Inversion [175], DPL [176], DiffusionDisentanglement [177],
Prompt Tuning Inversion [178] , leaf, text width=22.4em ] ] [ Guidance with
a Hypernetwork, text width=5em [ StyleDiffusion [179], InST [180] , leaf, text width=10.5em ] ] [ Latent Variable
Optimization, text width=5em [ DragonDiffusion [181], DragDiffusion [182], DDS [183],
DiffuseIT [184], CDS [185], MagicRemover [186] , leaf, text width=18em ] ] [ Hybrid
Finetuning, text width=3.5em [ Imagic [187], LayerDiffusion [188], Forgedit [189], SINE [190] , leaf, text width=20em ] ] ]
@@ 5.5 Hybrid Finetuning
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=2em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=4.5em,font=,, where level=2text width=3.1em,font=,, where level=3text width=3.0em,font=,, [ Training and Finetuning
Free Approaches [ Input Text
Refinement, text width=4em [ PRedItOR [191], ReDiffuser [192], Captioning and Injection [193], InstructEdit [194] , leaf, text width=26em ] ] [ Inversion/Sampling
Modification, text width=6em [ Direct Inversion [195], DDPM Inversion [196], SDE-Drag [197], LEDITS++ [198],
FEC [199], EMILIE [200], Negative Inversion [201], Proximal Inversion [202],
Null-Text Guidance [203], EDICT [204], AIDI [205], CycleDiffusion [206],
InjectFusion [207], Fixed-point inversion [208], TIC [209], Diffusion Brush [210],
Self-guidance [211] , leaf, text width=25.5em ] ] [ Attention
Modification, text width=4em [ P2P [212], Pix2Pix-Zero [213], MasaCtrl [85], PnP [214], TF-ICON [215],
Object-Shape Variations [216], Conditional Score Guidance [217], EBMs [218],
Shape-Guided Diffusion [219], HD-Painter [220] , leaf, text width=25em ] ] [ Mask
Guidance, text width=3em [ FISEdit [221], Blended Latent Diffusion [222], PFB-Diff [223], DiffEdit [224], RDM [225],
MFL [226], Differential Diffusion [227], Watch Your Steps [228], Blended Diffusion [229],
ZONE [230], Inpaint Anything [231] , leaf, text width=28em ] ] [ Multi-Noise
Redirection, text width=4em [ The Stable Artist [232], SEGA [233], LEDITS [234], OIR-Diffusion [235] , leaf, text width=23em ] ] ]
@@ 6 Training and Finetuning Free Approaches
In the field of image editing, training and finetuning free approaches start from the premise that they are fast and low-cost because they do not require any form of training (on the dataset) or finetuning (on the source image) during the complete process of editing. This section categorizes them into five categories according to what they modify, as shown in Figs. 7 and  8. They ingeniously leverage the inherent principles within the diffusion model for their editing goals.
@@ 6.4 Mask Guidance
Mask-Guided Regional Focus. This line of research employs masks as navigational tools, steering the sampling process towards specific regions of interest. The methods [227, 228] excel in localizing edits, ensuring that modifications align precisely with designated regions specified in text prompts or editing requirements. Differential Diffusion [227] pioneers a region-aware editing paradigm where masks are employed to customize the intensity of change within each region of the source image. This allows a granular level of control, fostering exquisite details and artistic expression. Watch Your Steps [228] utilizes masks as relevance maps to pinpoint desired edit regions, ensuring that only the most relevant regions are edited while preserving the other regions of the original image.
@@ 6.5 Multi-Noise Redirection
Object-Aware Noise Redirection. OIR-Diffusion [235] introduces the object-aware inversion and reassembly method, which enables fine-grained editing of specific objects within an image by determining optimal inversion steps for each sample and seamlessly integrating the edited regions with the other regions. It significantly maintains the fidelity of the non-edited regions and achieves remarkable results in editing diverse object attributes, particularly in intricate multi-object scenarios.
@@ 7.1.1 Supervised Training
[…]

### v3

@@ 3 Scope and Categorization
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=2em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, transform shape, , where level=1text width=4.5em,font=,, where level=2text width=3.1em,font=,, where level=3text width=3.0em,font=,, [ Training-Based
Approaches (Section 4 ) [ Domain-Specific Editing, text width=7.5em [ DiffusionCLIP [154], Asyrp [155], EffDiff [156], DiffStyler [157], StyleDiffusion [158], UNIT-DDPM [159],
CycleNet [160], Diffusion Autoencoders [161], HDAE [162], EGSDE [163], Pixel-Guided Diffusion [164] , leaf, text width=33em ] ] [ Reference and Attribute
Guided Editing, text width=7.5em [ PbE [165], RIC [166], ObjectStitch [167], PhD [168], DreamInpainter [169], Anydoor [170], FADING [171],
PAIR Diffusion [172], SmartBrush [173], IIR-Net [174], PowerPaint [175], Imagen Editor [176],
SmartMask [177],Uni-paint [178] , leaf, text width=33em ] ] [ Instructional Editing, text width=6.5em [ InstructPix2Pix [179], MoEController [180], FoI [181], LOFIE [182], InstructDiffusion [183], Emu Edit [184],
DialogPaint [185], Inst-Inpaint [186], HIVE [187], ImageBrush [188], InstructAny2Pix [189], MGIE [190],
SmartEdit [191] , leaf, text width=34em ] ] [ Pseudo-Target Retrieval Based Editing, text width=12em [ iEdit [192], TDIELR [193], ChatFace [194] , leaf, text width=14em ] ] ]
@@
Scope. In this survey, we concentrate on those works that most directly contribute to the field of image editing using diffusion models. To ensure a clear and focused discussion, this survey applies two primary criteria for the inclusion of works: (1) the task must focus on image editing, defined as modifying the appearance, structure, or content of existing images without significant transformations that would constitute new image generation, and (2) the methodology must depend on diffusion models performed on 2D images. Hence, works that utilize traditional image editing techniques or other generative models are excluded. Furthermore, works that rely on 3D knowledge are also not reviewed.
Categorization. In this survey, we organize diffusion model-based image editing papers into three principle groups based on their learning strategies: training-based approaches, testing-time finetuning approaches, and training and finetuning free approaches, which are elaborated in Sections 45, and 6, respectively. These learning strategies are defined by whether the methods require a substantial training phase, involve finetuning during the inference stage, or can perform without either, thus providing a comprehensive understanding of the resources and processes involved in applying these methods. Additionally, we explore 10 types of input conditions employed by these methods to control the editing process, including text, mask, reference (Ref.) image, class, layout, pose, sketch, segmentation (Seg.) map, audio, and dragging points. Furthermore, we investigate 12 most common editing types that can be accomplished by these methods, which are organized into three broad categories defined as follows.
@@ 4 Training-Based Approaches
In the field of diffusion model-based image editing, training-based approaches refer to methods that require a substantial training phase on large datasets before they can be applied to specific editing tasks. These methods are notable for their stable training processes and effective modeling of data distribution, leading to reliable performance across a variety of editing scenarios. This section categorizes these methods into four main groups based on their application scopes, and the conditions required for training, as shown in Fig. 2. Further, within each of these primary groups, we classify the approaches into distinct types based on their core editing methodologies. This classification illustrates the range of these methods, from targeted domain-specific applications to broader open-world uses.
4.1 Domain-Specific Editing
Over the past several years, Generative Adversarial Networks (GANs) have been widely explored in image editing for their ability to generate high-quality images. However, diffusion models, with their advanced image generation capabilities, emerge as a new focus in this field. One challenge with diffusion models is their need for extensive computational resources when trained on large datasets. To address this, earlier studies train these models on smaller specialized datasets, which highly concentrate on specific domains such as CelebA [262] and FFHQ [11] for human face manipulation, AFHQ [263] for animal face editing and translation, LSUN [264] for object modification, and WikiArt [265] for style transfer. To thoroughly understand these approaches, we organize them according to their types of supervision.
CLIP Guidance. Drawing inspiration from GAN-based methods [266, 267] that use CLIP [268] for guiding image editing with text, several studies incorporate CLIP into diffusion models. A key example is DiffusionCLIP [154], which allows for image manipulation in both trained and new domains using CLIP. Specifically, it first converts a real image into latent noise using DDIM inversion, and then finetunes the pretrained diffusion model during the reverse diffusion process to adjust the image’s attributes constrained by a CLIP loss between the source and target text prompts. Instead of finetuning on the whole diffusion model, Asyrp [155] focuses on a semantic latent space internally, termed h-space, where it defines an additional implicit function parameterized by a small neural network. Then it trains the network with the guidance of the CLIP loss while keeping the diffusion model frozen. To address the time-consuming problem of multi-step optimization in DiffusionCLIP, EffDiff [156] introduces a faster method with single-step training and efficient processing.
@@ 4.1 Domain-Specific Editing
Cycling Regularization. Since diffusion models are capable of domain translation, the cycling framework used in methods like CycleGAN [14] is also explored within them. For example, UNIT-DDPM [159] defines a dual-domain Markov chain within diffusion models, using cycle consistency to regularize training for unpaired image-to-image translation. Similarly, CycleNet [160] adopts ControlNet [102] with pretrained Stable Diffusion [30] as the backbone for text conditioning. It employs consistency regularization throughout the image translation cycle, encompassing both forward translation from the source domain to the target domain and backward translation in the reverse direction.
@@
4.2 Reference and Attribute Guided Editing
@@
4.3 Instructional Editing
@@
4.4 Pseudo-Target Retrieval Based Editing
@@ 5 Testing-Time Finetuning Approaches
Testing-time finetuning approaches usually do not involve a traditional pre-training phase but instead require finetuning during the inference (testing-time) stage for each specific image to be edited. These methods are designed to optimize the model’s performance for individual images, allowing for more precise and personalized edits. This section explores various finetuning strategies, as shown in Fig. 4 and Fig.  5, ranging from finetuning the entire model to adjusting specific layers or embeddings. These methods collectively demonstrate the evolving complexity and effectiveness of finetuning techniques in image editing, catering to various editing requirements and user intentions.
@@ 5.1 Denoising Model Finetuning
Finetuning Entire Denoising Models. Finetuning the entire denoising model allows the model to better learn specific features of images and more accurately interpret textual prompts, resulting in edits that more closely align with user intent. UniTune [195] finetunes the diffusion model on a single base image during the tuning phase, encouraging the model to produce images similar to the base image. During the sampling phase, a modified sampling process is used to balance fidelity to the base image and alignment to the editing prompt. This includes starting sampling from a noisy version of the base image and applying classifier-free guidance during the sampling process. Custom-Edit [196] uses a small set of reference images to personalize the diffusion model, enhancing the similarity of the edited image to the reference images while maintaining the similarity to the source image.
@@
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=2em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, transform shape, , where level=1text width=4.5em,font=,, where level=2text width=3.1em,font=,, where level=3text width=3.0em,font=,, [ Testing-Time Finetuning
Approaches (Section  5) [ Denoising Model Finetuning, text width=9em [ UniTune [195], Custom-Edit [196], KV-Inversion [197] , leaf, text width=17em ] ] [ Embedding Finetuning, text width=7.5em [ Null-Text Inversion [198], DPL [199], DiffusionDisentanglement [200], Prompt Tuning Inversion [201] , leaf, text width=32em ] ] [ Guidance with a Hypernetwork, text width=10em [ StyleDiffusion [202], InST [203] , leaf, text width=10.5em ] ] [ Latent Variable Optimization, text width=9em [ DragonDiffusion [204], DragDiffusion [205], DDS [206], DiffuseIT [207], CDS [208], MagicRemover [209] , leaf, text width=33em ] ] [ Hybrid Finetuning, text width=6em [ Imagic [210], LayerDiffusion [211], Forgedit [212], SINE [213] , leaf, text width=20em ] ] ]
@@ 5.4 Latent Variable Optimization
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=2em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, transform shape, , where level=1text width=4.5em,font=,, where level=2text width=3.1em,font=,, where level=3text width=3.0em,font=,, [ Training and Finetuning Free
Approaches (Section 6) [ Input Text Refinement, text width=7em [ PRedItOR [214], ReDiffuser [215], Captioning and Injection [216], InstructEdit [217] , leaf, text width=26em ] ] [ Inversion/Sampling
Modification, text width=6em [ Direct Inversion [218], DDPM Inversion [219], SDE-Drag [220], LEDITS++ [221], FEC [222],
EMILIE [223], Negative Inversion [224], Proximal Inversion [225], Null-Text Guidance [226],
EDICT [227], AIDI [228], CycleDiffusion [229], InjectFusion [230], Fixed-point inversion [231],
TIC [232], Diffusion Brush [233], Self-guidance [234] , leaf, text width=30em ] ] [ Attention
Modification, text width=4em [ P2P [235], Pix2Pix-Zero [236], MasaCtrl [87], PnP [237], TF-ICON [238], Object-Shape Variations [239],
Conditional Score Guidance [240], EBMs [241], Shape-Guided Diffusion [242], HD-Painter [243] , leaf, text width=33em ] ] [ Mask
Guidance, text width=3em [ FISEdit [244], Blended Latent Diffusion [245], PFB-Diff [246], DiffEdit [247], RDM [248], MFL [249],
Differential Diffusion [250], Watch Your Steps [251], Blended Diffusion [252], ZONE [253], Inpaint Anything [254] , leaf, text width=36em ] ] [ Multi-Noise Redirection, text width=7.5em [ The Stable Artist [255], SEGA [256], LEDITS [257], OIR-Diffusion [258] , leaf, text width=23em ] ] ]
@@ 6 Training and Finetuning Free Approaches
Training and finetuning free approaches start from the premise that they are fast and low-cost. They do not require any form of training (on the dataset) or finetuning (on the source image) during the editing process. This section categorizes these methods into five groups based on the specific elements they modify within the image, as shown in Figs. 6 and 7, highlighting how they leverage the inherent principles of diffusion models for efficient editing.
@@ 6.4 Mask Guidance
Mask-Guided Regional Focus. This line of research employs masks as navigational tools, steering the sampling process towards specific regions of interest. The methods [250, 251] excel in localizing edits, ensuring that modifications align precisely with designated regions specified in text prompts or editing requirements. Differential Diffusion [250] pioneers a region-aware editing paradigm where masks are employed to personalize the intensity of change within each region of the source image. This allows a granular level of control, fostering exquisite details and artistic expression. Watch Your Steps [251] utilizes masks as relevance maps to pinpoint desired edit regions, ensuring that only the most relevant regions are edited while preserving the other regions of the original image.
@@ 6.5 Multi-Noise Redirection
Object-Aware Noise Redirection.
OIR-Diffusion [258] introduces the object-aware inversion and reassembly method, which enables fine-grained editing of specific objects within an image by determining optimal inversion steps for each sample and seamlessly integrating the edited regions with the other regions. It significantly maintains the fidelity of the non-edited regions and achieves remarkable results in editing diverse object attributes, particularly in intricate multi-object scenarios.
@@ 7.1.1 Supervised Training
The supervised training-based inpainting & outpainting methods usually train a diffusion model from scratch with paired corrupted and complete images. Palette [45] develops a unified framework for image-to-image translation based on conditional diffusion models and individually trains this framework on four image-to-image translation tasks, namely colorization, inpainting, outpainting, and JPEG restoration. It utilizes a direct concatenation of low-quality reference images with the denoised result at the $t-1$ step as the condition for noise prediction at the $t$ step. SUD2 [290] builds upon and extends the SUD framework [291], integrating additional strategies like correlation minimization, noise injection, and particularly the use of denoising diffusion models to address the scarcity of paired training data in semi-supervised learning.